OpenStudy (anonymous):

What is the derivative of 4pi and why?

OpenStudy (anonymous):

0, because it's a constant.

OpenStudy (anonymous):

I thought it was 2pi

OpenStudy (anonymous):

Why do you think it'd be \(2\pi\)? Explain how you got to the conclusion.

OpenStudy (anonymous):

I am working on this problem right now dealing with Area. This is what it looks like: \[c^2\div2\pi \] and the answer books says the derivative is C/2pi

OpenStudy (anonymous):

wait.....C^2/4pi

OpenStudy (anonymous):

That's an entirely different problem. You actually have a variable, whose derivative is \(2C\), multiplying that by \(\frac{1}{4\pi}\) gives you: \[ \frac{2C}{4\pi}=\frac{C}{2\pi} \]. So, to answer your original question, yes: \[ \frac{d}{dC}4\pi=0 \]BUT: \[ \frac{d}{dC}\frac{C^2}{4\pi}=\frac{2C}{4\pi}=\frac{C}{2\pi} \]Which is different.

OpenStudy (anonymous):

always remember that the derivative of constant function is 0 ( zero ) .. 4pi is constant function...